# Context is Key: A Benchmark for Forecasting with Essential Textual Information

### v1

@@ C.3 Significant failures per model
     Time-LLM 2
C.4 Visualizations of successful context-aware forecasts
@@
Context: “ This series contains Diffuse Horizontal Irradiance for a location in Sinaloa, Mexico. The Diffuse Horizontal Irradiance is the total amount of sun energy (in Watts per squared meter) arriving indirectly on a horizontal surface, ignoring the direct sunlight. Even when there are no clouds to scatter the sun light, there will still be some Diffuse Horizontal Irradiance, since clouds are not the only cause of light scattering. When there are no clouds, the Diffuse Horizontal Irradiance is mostly a function of the position of the sun in the sky, with only small variations from factors such as water vapour and dust particles levels. If the cloud cover is light, the Diffuse Horizontal Irradiance will increase due to the increase scattering of sun light, but heavy cloud cover will decrease it due to some sun light no longer being able to reach the ground. At the beginning of the series, the weather was cloudy.
At 2022-07-12 11:00:00, the weather became clear.
At 2022-07-12 19:00:00, the weather became cloudy.
At 2022-07-13 12:00:00, the weather became clear.
At 2022-07-13 13:00:00, the weather became cloudy.
At 2022-07-14 06:00:00, we expect that the weather will become clear.
At 2022-07-14 07:00:00, we expect that the weather will become cloudy.
At 2022-07-14 10:00:00, we expect that the weather will become clear.
At 2022-07-14 18:00:00, we expect that the weather will become cloudy. ”
@@
Context: “ This is the Unemployment Rate for Okaloosa County, in Florida. For reference, here is the Unemployment Rate for a few American states during the same period:
Pennsylvania
@@
Figure 14: Example of successful context-aware forecasts by LLMP with Llama-70B
@@
Figure 15: Example of successful context-aware forecasts by LLMP with Llama-70B
C.5 Visualizations of significant failures
Context: “ Given are variables X_0 and X_1, where X_0 is a covariate and X_1 is the variable to forecast. Variables are generated from a linear Structural Vector Autoregressive (SVAR) model with additive gauss noise and a noise scale of 1.487e-03, with lag = 3. The task is to forecast the value of the variable X_1 at time t, given the values of the covariate X_0 and the variable X_1 itself at times t-1, … t-3. For the first 128 days, the covariate X_0 takes a value of 8 from 2024-02-21 to 2024-03-11, 12 from 2024-03-12 to 2024-05-06, 12 from 2024-05-07 to 2024-06-27. For the next 32 days, the covariate X_0 takes a value of 30 from 2024-06-28 to 2024-07-13, 60 from 2024-07-14 to 2024-07-14, 60 from 2024-07-15 to 2024-07-29. Each day can be treated as a timestep for the forecasting task. The causal parents affect the child variables at different lags. The causal parents for each variable is given below:
No parents for X_0 at any lag.
Parents for X_1 at lag 1: [’X_0’, ’X_1’] affect the forecast variable as 0.527 * X_0 + -0.895 * X_1.
Parents for X_1 at lag 2: [’X_0’, ’X_1’] affect the forecast variable as 1.380 * X_0 + -0.758 * X_1.
Parents for X_1 at lag 3: [’X_0’, ’X_1’] affect the forecast variable as -0.661 * X_0 + -0.793 * X_1. ”
@@ C.5 Visualizations of significant failures
C.6 Cost of API-based models
Model Total Per-instance average Per-instance median
GPT-4o \$143.83 \$0.288 \$0.170
GPT-4o (no context) \$139.50 \$0.279 \$0.160
GPT-4o-mini \$13.79 \$0.040 \$0.040
GPT-4o-mini (no context) \$13.32 \$0.038 \$0.040
C.7 Impact of Irrelevant Information in Context
As a way to get some insight on this question, we took a task where the models have to forecast the unemployment data of an American county, given the unemployment data of the state the county is in. We then modify this task by first trying to mislead the model by wrongly saying that the state-level data was from another state (without changing the data itself), then by giving the data from the other state (while explicitly telling the model that data is from said other state), before finally removing the state-level data altogether. The result for this experiment with 5 instances per task for Direct Prompt - GPT-4o is shown in Tab. 6, while the forecasts for a single instances are shown in Fig. 20. From these, we see that the model aggressively used data which is marked as being from an other state, even though if the data was actually from said other state, the performance would be closer to not having any state-level data. This shows that the model is liable to take any information provided as being useful, even though its usefulness is marginal.
Table 6: Ability of the Direct Prompt - GPT-4o model to accurately predict the unemployement level of an American county, given various covariates. These results are averaged over 5 instances.
Available data RCPRS
Data from the correct state, accurately tagged 0.0583
Data from the correct state, inaccurately tagged 0.0557
Data from an incorrect state, accurately tagged 0.1966
No state-level data 0.2630
Figure 20: Forecasts done by Direct Prompt - GPT-4o, with varying information in the context. The task is to forecast the forecast the unemployment rate of an American county.
@@ D.1.1 Method
We observe that models often produce samples that fail to adhere to the structure and are therefore rejected. When sampling $25$ samples from the model, with Direct Prompt, we allow retrying for a maximum of $K$ times until we obtain $25$ valid samples. If we do not have $25$ valid samples from the model at the end of $K$ retries, we record a failure of the model and attribute the model the RCRPS upper bound of 5 for that task. In practice, we find that larger models (Llama-3.1-405B-Instruct, GPT-4o and GPT-4o) can produce $25$ valid forecasts with $1$ to $3$ retries. However with the smaller models (such as Llama-3-70B-Instruct, Llama-3-8B-Instruct and Mixtral-8x7B-Instruct), up to $10$ retries can be required to obtain $25$ valid forecasts. Further, we found that without an explicit “Do not include any other information (e.g., comments) in the forecast.”, models often included unwanted information along with the forecasts.
Instruction-tuning is necessary for models to work with Direct Prompt
Direct Prompting requires forecasts to be produced in a specific structure. To generate structured outputs, models need to be steerable (Dubey et al., 2024), a capability that is typically elicited from base models with post-training methods such as instruction tuning (Wei et al., 2021). We observe this in our evaluations as we find that several base models, including Llama-3-8B, Llama-3-70B, Mixtral-8x7B, and even the biggest base model we tried, Llama-3.1-405B, are incapable of generating outputs adhering to the structure required for Direct Prompt, despite increasing the number of retries to as high as $50$ retries. With Direct Prompt, these models often output irrelevant information, sometimes completing solely the context as a text completion task, and in other cases regurgitating forecasting datasets that they have memorized.
@@ D.1.2 Implementation Details
We used a single H100 GPU to run the Direct Prompt approach for Llama-3-8B-Instruct, and 2 H100 GPUs for Llama-3-70B-Instruct and Mixtral-8x7B-Instruct. We queried Llama-3.1-405b-Instruct from an externally-hosted server running on 8 H100s. We use the OpenAI API to perform inference on the proprietary GPT-4o and GPT-4o-mini models. We provide the cost incurred in the inference of these models with the Direct Prompt method in Sec. C.6.
@@ D.2 LLMP
We used a single H100 GPU to run the LLMP approach for the following models: Llama-3-8B, and Llama-3-8B-Instruct. We used 2 H100 GPUs for the following models: Mixtral-8x7B, and Mixtral-8x7B-Instruct, and used used 8 H100 GPUs for the following models: Llama-3-70B, and Llama-3-70B-Instruct.
@@ D.3 UniTime and Time-LLM
For multimodal models, we jointly train UniTime (Liu et al., 2024b) on its ensemble of datasets: ETTm1, ETTm2, ETTh1, ETTh2, Electricity, Weather, Exchange and Illness.
@@
UniTime:  We train UniTime (Liu et al., 2024b) using a single seed on one AMD Instinct MI200 GPU for approximately 14 hours. It features a lightweight transformer with maximum context length of 210 and a pre-trained GPT2 language model as backbone, of which only the first half of the transformer layers are used. The time series baseline employs non-overlapping patch embeddings generated with a kernel size and stride of 16, and a maximum input sequence length of 96. When the total tokenized length exceeds the architecture’s capacity, we truncate the context.
@@
Time-LLM:  We train Time-LLM (Jin et al., 2024) on the ETTh1 dataset (Zhou et al., 2021) with a prediction length of 96. We train using a single seed on four AMD Instinct MI200 GPUs, with an average training time per run of approximately 13 hours. Training was conducted using a batch size of 8 per device and 4 gradient accumulation steps, along with a 1Cycle learning rate schedule with a maximum learning rate of 1e-3. In addition, runs were accelerated using DeepSpeed Stage 2 and BF16 automatic mixed precision.
@@
Interestingly, UniTime’s performance does benefit from context for some tasks (see Fig. 21). However, the aggregate RCRPS and rank of UniTime with respect to other models indicate that it still struggles to produce forecasts competitive with even quantitative forecasting methods.

### v2

@@ C.5 Significant failures per model
     Time-LLM (ETTh1) 2
Without Context
direct prompt
Llama-3.1-405B-Inst
9
Llama-3-70B-Inst
4
Llama-3-8B-Inst
8
Mixtral-8x7B-Inst
14
Qwen-2.5-7B-Inst
0
Qwen-2.5-1.5B-Inst
5
Qwen-2.5-0.5B-Inst
0
GPT-4o
2
GPT-4o-mini
1
llmp
Llama-3-70B-Inst
1
Llama-3-70B
0
Llama-3-8B-Inst
9
Llama-3-8B
3
Mixtral-8x7B-Inst
3
Mixtral-8x7B
0
Qwen-2.5-7B-Inst
35
Qwen-2.5-7B
16
Qwen-2.5-1.5B-Inst
57
Qwen-2.5-1.5B
37
Qwen-2.5-0.5B-Inst
45
Qwen-2.5-0.5B
74
Multimodal Models
UniTime
1
Time-LLM (ETTh1)
1
TS Foundation Models
Lag-Llama
1
Chronos-Tiny
2
Chronos-Mini
2
Chronos-Small
1
Chronos-Base
18
Chronos-Large
1
TimeGEN
2
Moirai-Small
3
Moirai-Base
8
Moirai-Large
7
Statistical Models
ARIMA
2
ETS
1
Exp-Smoothing
4
We observe that in a few instances in the benchmark, some models tend to obtain significantly worse performance when evaluated with context. In our evaluation, we term all instances where the RCRPS value of a model is greater than $5$, as significant failures of the model on those instances. We found $5$ as a suitable value for analyzing such failures, as it intuitively represents the value a forecast would get if the distance between the forecast and the ground-truth was 5 times bigger than the range of the ground-truth for the task. When we aggregate the RCRPS of instances in the benchmark (such as in Tab. 1), we cap the RCRPS of such significant failures to $5$, to avoid outliers with a much higher RCRPS affecting the aggregate score. In Tab. 5, we show the number of such instances in our evaluation of the benchmark where we found models to have significant failures (out of a total of $355$ evaluated instances). Interestingly, some models such as direct prompt with Llama-3.1-405B-Instruct and llmp with Llama-3-70B and Llama-3-8B are more robust to such significant failures, and do not incur such failures. On the other hand, models such as Qwen family of models (that are notably significantly smaller than the rest) with llmp achieve the most significant failures, followed by Llama-3-70B-Instruct and Llama-3-8B-Instruct with llmp. We postulate that this is because of models misinterpreting context. It is still an open question as to how to increase the robustness of models to prevent or reduce such significant failures. We visualize such significant failures in Sec. C.12.
C.6 Testing the Statistical Significance of the Relevance of Context
Table 6: $p$-value of the one-sided paired $t$-test between the RCPRS values with and without context for models who can use it. Since this test is done on the unweighted RCRPS values, the average RCPRS presented in this table are also unweighted.
Average RCRPS Average RCRPS
Model With context Without context $p$-value
direct prompt
Llama-3.1-405B-Inst
0.165 ± 0.005 0.544 ± 0.007 7.54 $\times$ 10-${}^{\text{-13}}$
llmp -
Llama-3-70B
0.192 ± 0.004 0.249 ± 0.004 2.09 $\times$ 10-${}^{\text{-9}}$
llmp -
Mixtral-8x7B
0.202 ± 0.005 0.245 ± 0.004 9.17 $\times$ 10-${}^{\text{-8}}$
llmp -
Llama-3-8B
0.216 ± 0.007 0.283 ± 0.003 5.95 $\times$ 10-${}^{\text{-4}}$
llmp -
Mixtral-8x7B-Inst
0.223 ± 0.002 0.290 ± 0.009 0.002
direct prompt
Qwen-2.5-7B-Inst
0.243 ± 0.003 0.402 ± 0.009 8.01 $\times$ 10-${}^{\text{-8}}$
direct prompt
Llama-3-70B-Inst
0.246 ± 0.003 0.529 ± 0.010 1.07 $\times$ 10-${}^{\text{-10}}$
direct prompt
GPT-4o-mini
0.249 ± 0.003 0.403 ± 0.005 2.13 $\times$ 10-${}^{\text{-8}}$
direct prompt
GPT-4o
0.254 ± 0.010 0.387 ± 0.007 7.44 $\times$ 10-${}^{\text{-4}}$
UniTime
0.290 ± 0.001 0.321 ± 0.001 0.016
direct prompt
Qwen-2.5-0.5B-Inst
0.344 ± 0.011 0.258 ± 0.004 0.988
Time-LLM (ETTh1)
0.378 ± 0.001 0.364 ± 0.001 1.000
llmp -
Llama-3-8B-Inst
0.413 ± 0.009 0.432 ± 0.010 0.294
direct prompt
Mixtral-8x7B-Inst
0.414 ± 0.007 0.699 ± 0.006 1.89 $\times$ 10-${}^{\text{-6}}$
direct prompt
Llama-3-8B-Inst
0.415 ± 0.007 0.631 ± 0.007 3.42 $\times$ 10-${}^{\text{-10}}$
direct prompt
Qwen-2.5-1.5B-Inst
0.480 ± 0.016 0.406 ± 0.006 0.973
llmp -
Llama-3-70B-Inst
0.591 ± 0.019 0.313 ± 0.003 1.000
llmp -
Qwen-2.5-7B
0.902 ± 0.025 0.619 ± 0.018 1.000
llmp -
Qwen-2.5-1.5B-Inst
2.031 ± 0.025 1.180 ± 0.022 1.000
llmp -
Qwen-2.5-0.5B-Inst
2.067 ± 0.025 1.046 ± 0.017 1.000
llmp -
Qwen-2.5-0.5B
2.139 ± 0.020 1.765 ± 0.025 1.000
llmp -
Qwen-2.5-1.5B
2.162 ± 0.028 0.861 ± 0.016 1.000
llmp -
Qwen-2.5-7B-Inst
2.282 ± 0.025 0.899 ± 0.014 1.000
To assess whether the lower average RCPRS using context than without using context we observe for the best performing model in our benchmark is statistically significant, we ran an analysis using the paired $t$-test. We used the paired $t$-test implemented in the scipy Python package as the scipy.stats.ttest_rel method, with alternative="less" to make the test one-sided. As can be seen in Tab. 6, for many models the improved RCPRS when using the context is statistically significant, with $p$-values lower than $\times$ 10-${}^{\text{-6}}$ for 7 out of the 23 models under consideration. Furthermore, the best performing models are those for which the improvement is statistically significant, with the 9 best models all having $p$-values below 0.01.
C.7 Cost of API-based models
Model Total Per-instance average Per-instance median
GPT-4o
\$143.83 \$0.288 \$0.170
GPT-4o
(no context)
\$139.50 \$0.279 \$0.160
GPT-4o-mini
\$13.79 \$0.040 \$0.040
GPT-4o-mini
(no context)
\$13.32 \$0.038 \$0.040
C.8 Impact of Relevant and Irrelevant Information in Context
Figure 14: A comparison of RCRPS (lower is better) for two tasks on predicting the Unemployment Rate of a county. Both contain the context needed to solve the task. However, the UnemploymentCountyUsingSingleStateData task (dark green) is filtered to only contain the relevant context. Other the other hand, the UnemploymentCountyUsingExpliciteMultipleStateData task (light green) also contains other unrelated context. We visualize three models here, all of which perform better when the context only includes the most relevant information.
We study here if models perform better on context that has already been filtered to only contain relevant information. To assess this, we compare two tasks on predicting the Unemployment Rate of a county.
  1. 1.
    For the UnemploymentCountyUsingSingleStateData task, the context contains the unemployment rate of the state which the county belongs to, tagged with the name of the state. See https://anon-forecast.github.io/benchmark_report_dev/UnemploymentCountyUsingSingleStateData.html for a visualization.
  2. 2.
[…]
Results of three randomly picked models from the benchmark is visualized in Figure 14. We find that models perform much better when only the relevant state’s data is provided, as opposed to the context also containing data from other states.
C.9 Impact of Solely Irrelevant Information in Context
[…]
Table 8: Ability of the direct prompt  - GPT-4o model to accurately predict the unemployment level of an American county, given various covariates. These results are averaged over 5 instances.
Available data RCPRS
Data from the correct state, accurately tagged 0.0583
Data from the correct state, inaccurately tagged 0.0557
Data from an incorrect state, accurately tagged 0.1966
No state-level data 0.2630
Figure 15: Forecasts done by Direct Prompt - GPT-4o, with varying information in the context. The task is to forecast the unemployment rate of an American county.
C.10 The effect of significant failures on the aggregate performance of models
Figure 16: Histogram of the RCPRS (lower is better) of the direct prompt  - Mixtral-8x7B-Instruct model on each task, with and without context (with the weighting scheme detailed in Sec. A.5). With context, the model gets an RCRPS close to zero in a large number of tasks (also achieving a high win rate as seen in Figure 5). However, there is also a long tail of tasks with high RCRPS values with context, dominating and worsening the model’s aggregate RCRPS.
As discussed in Sec. 5.4, in a few instances from the benchmark, some models return forecasts that miss the ground truth by a large margin, which we term significant failures (detailed in Sec. C.5). We analyse the effect of such significant failures on the results here. We use the direct prompt  - Mixtral 8x7B model as an example here, while the same phenomenon may apply to other models. In Figure 6, we can find that the aggregate RCRPS of direct prompt  - Mixtral 8x7B worsens when it uses context. However, in Figure 5 (left), the win rate of the model vs quantitative baselines improves when it uses context. These two figures show results that seem contradictory, but are in fact compatible: adding context improves the model’s RCRPS for most tasks, but greatly worsens it for a minority of tasks where the model achieves significant failures.
To further illustrate this effect, we visualize the task-wise RCRPS of the direct prompt  - Mixtral-8x7B-Instruct model, both with and without context, in Figure 16. With context, the model gets an RCRPS close to zero in a large number of tasks. However, there is also a long tail of tasks with high RCRPS values with context, dominating and worsening the model’s aggregate RCRPS.
C.11 Visualizations of successful context-aware forecasts
@@
Context: “ This series contains Diffuse Horizontal Irradiance for a location in Sinaloa, Mexico. The Diffuse Horizontal Irradiance is the total amount of sun energy (in Watts per squared meter) arriving indirectly on a horizontal surface, ignoring the direct sunlight. Even when there are no clouds to scatter the sun light, there will still be some Diffuse Horizontal Irradiance, since clouds are not the only cause of light scattering. When there are no clouds, the Diffuse Horizontal Irradiance is mostly a function of the position of the sun in the sky, with only small variations from factors such as water vapour and dust particles levels. If the cloud cover is light, the Diffuse Horizontal Irradiance will increase due to the increase scattering of sun light, but heavy cloud cover will decrease it due to some sun light no longer being able to reach the ground. At the beginning of the series, the weather was cloudy. At 2022-07-12 11:00:00, the weather became clear. At 2022-07-12 19:00:00, the weather became cloudy. At 2022-07-13 12:00:00, the weather became clear. At 2022-07-13 13:00:00, the weather became cloudy. At 2022-07-14 06:00:00, we expect that the weather will become clear. At 2022-07-14 07:00:00, we expect that the weather will become cloudy. At 2022-07-14 10:00:00, we expect that the weather will become clear. At 2022-07-14 18:00:00, we expect that the weather will become cloudy. ”
@@
Context: “ This is the Unemployment Rate for Okaloosa County, in Florida. For reference, here is the Unemployment Rate for a few American states during the same period: Pennsylvania
@@
Figure 23: Example of successful context-aware forecasts by llmp with Llama-3-70B
@@
Figure 24: Example of successful context-aware forecasts by llmp with Llama-3-70B
C.12 Visualizations of significant failures
Context: “ Given are variables X_0 and X_1, where X_0 is a covariate and X_1 is the variable to forecast. Variables are generated from a linear Structural Vector Autoregressive (SVAR) model with additive gauss noise and a noise scale of 1.487e-03, with lag = 3. The task is to forecast the value of the variable X_1 at time t, given the values of the covariate X_0 and the variable X_1 itself at times t-1, … t-3. For the first 128 days, the covariate X_0 takes a value of 8 from 2024-02-21 to 2024-03-11, 12 from 2024-03-12 to 2024-05-06, 12 from 2024-05-07 to 2024-06-27. For the next 32 days, the covariate X_0 takes a value of 30 from 2024-06-28 to 2024-07-13, 60 from 2024-07-14 to 2024-07-14, 60 from 2024-07-15 to 2024-07-29. Each day can be treated as a timestep for the forecasting task. The causal parents affect the child variables at different lags. The causal parents for each variable is given below: No parents for X_0 at any lag. Parents for X_1 at lag 1: [’X_0’, ’X_1’] affect the forecast variable as 0.527 * X_0 + -0.895 * X_1. Parents for X_1 at lag 2: [’X_0’, ’X_1’] affect the forecast variable as 1.380 * X_0 + -0.758 * X_1. Parents for X_1 at lag 3: [’X_0’, ’X_1’] affect the forecast variable as -0.661 * X_0 + -0.793 * X_1. ”
@@ D.1.1 Method
To constrain the output of the model to follow the specific structure, we use constrained decoding using the lm-format-enforcer tool (https://github.com/noamgat/lm-format-enforcer) and a regular expression that only allows models to output the values corresponding to the prediction timestamps. Without constrained decoding, we observe that models often produce samples that fail to adhere to the structure and are therefore rejected. Specifically, larger models ( Llama-3.1-405B-Instruct , GPT-4o and GPT-4o-mini ) can produce $25$ valid forecasts with $1$ to $3$ trials. However with the smaller models (such as Llama-3-70B-Instruct , Llama-3-8B-Instruct and Mixtral-8x7B-Instruct ), up to $10$ trials can be required to obtain $25$ valid forecasts.
Further, we found that without an explicit “Do not include any other information (e.g., comments) in the forecast.”, models often included unwanted information along with the forecasts.
Instruction-tuned models are more amenable for direct prompt
direct prompt
requires forecasts to be produced in a specific structure. To generate structured outputs, models need to be steerable (Dubey et al., 2024), a capability that is typically elicited from base models with post-training methods such as instruction tuning (Wei et al., 2021). We observe this in our evaluations as we find that several base models, including Llama-3-8B , Llama-3-70B , Mixtral-8x7B , and even the biggest base model we tried, Llama-3.1-405B , are incapable of generating outputs adhering to the structure required for direct prompt , despite increasing the number of retries to as high as $50$ retries. With direct prompt , these models often output irrelevant information, sometimes completing solely the context as a text completion task, and in other cases regurgitating forecasting datasets that they have memorized.
@@ D.1.2 Implementation Details
We used a single H100 GPU to run the direct prompt approach for Llama-3-8B-Instruct , and 2 H100 GPUs for Qwen-2.5-{0.5B-Instruct, 1.5-Instruct, 7B-Instruct} , Llama-3-70B-Instruct and Mixtral-8x7B-Instruct . We queried Llama-3.1-405b-Instruct from an externally-hosted server running on 8 H100s. We use the OpenAI API to perform inference on the proprietary GPT-4o and GPT-4o-mini models. We provide the cost incurred in the inference of these models with the direct prompt method in Sec. C.7.
@@ D.2.1 Method
D.2.2 Implementation Details
We used a single H100 GPU to run the llmp approach for the following models: Llama-3-8B , and Llama-3-8B-Instruct . We used 2 H100 GPUs for the Qwen-2.5 family of models, Mixtral-8x7B , and Mixtral-8x7B-Instruct , and used used 8 H100 GPUs for the following models: Llama-3-70B , and Llama-3-70B-Instruct .
Since the code of llmp (https://github.com/requeima/llm_processes/) only supports using open-source models (such as those available in HuggingFace) and requires loading the weights into memory, it does not support experimenting with the GPT-4o and GPT-4o-mini models. Further, due to the memory requirements of llmp, we were unable to experiment with the Llama-3.1-405B and Llama-3.1-405B-Instruct models that required more than 24 H100 GPUs in parallel to process a single instance from the benchmark, which exceeded our available resources.
@@ D.3 UniTime and Time-LLM
For multimodal models, we jointly train UniTime (Liu et al., 2024c) on its ensemble of datasets: ETTm1, ETTm2, ETTh1, ETTh2, Electricity, Weather, Exchange and Illness.
@@
UniTime
:  We train UniTime  (Liu et al., 2024c) with their codebase (https://github.com/liuxu77/UniTime) using a single seed on one AMD Instinct MI200 GPU for approximately 14 hours. It features a lightweight transformer with maximum context length of 210 and a pre-trained GPT2 language model as backbone, of which only the first half of the transformer layers are used. The time series baseline employs non-overlapping patch embeddings generated with a kernel size and stride of 16, and a maximum input sequence length of 96. When the total tokenized length exceeds the architecture’s capacity, we truncate the context.
@@
Time-LLM
:  We train Time-LLM  (Jin et al., 2024) with their codebase (https://github.com/KimMeen/Time-LLM) on the ETTh1 dataset (Zhou et al., 2021) with a prediction length of 96. We train using a single seed on four AMD Instinct MI200 GPUs, with an average training time per run of approximately 13 hours. Training was conducted using a batch size of 8 per device and 4 gradient accumulation steps, along with a 1Cycle learning rate schedule with a maximum learning rate of 1e-3. In addition, runs were accelerated using DeepSpeed Stage 2 and BF16 automatic mixed precision.
@@
Interestingly, UniTime ’s performance does benefit from context for some tasks (see Figure 29). However, the aggregate RCRPS and rank of UniTime with respect to other models indicate that it still struggles to produce forecasts competitive with even quantitative forecasting methods.